OpenStudy (anonymous):

Integration by partial fractions... integrate (5x+2) / (2x^2 + x) Pleaseeee show your steps, I keep getting: 2ln(x) + ln(2x + 1) + C. But I believe this is not the correct answer.

OpenStudy (anonymous):

wolfram says answer is 2ln(x) + (1/2)ln(2x + 1) + C i didn't try to solve it but do you see where the mistake is?

OpenStudy (tkhunny):

It's REALLY close. Wolfram does not quite have it right. Should be: \(2ln(|x|) + (1/2)ln(|2x+1|) + C\) We don't know anything about x in the indefinite integral. The absolute values are NOT optional. Why is it integration by parts? Try partial fractions.

OpenStudy (anonymous):

My mistake I meant to say partial fractions... :P

OpenStudy (anonymous):

I don't see why there is a 1/2 on the second expression

OpenStudy (tkhunny):

\(\int \dfrac{1}{x+1}\;dx = ln(|x+1| + C)\) \(\int \dfrac{1}{2x+1}\;dx = (1/2)ln(|2x+1| + C)\) \(\int \dfrac{1}{\pi x+1}\;dx = (1/\pi)ln(|\pi x+1| + C)\) Think "Chain Rule".

OpenStudy (anonymous):

Wow I understand now... I skipped the substitution step and just put the expression in ln(_)

OpenStudy (anonymous):

but with substitution there will be a fraction, correct? in this case u = 2x + 1, du = 2 dx, dx = 1/2 du which is where the 1/2 comes from

OpenStudy (anonymous):

yes! thank you!

OpenStudy (tkhunny):

Well, there you have it. Good work.